the zero coupon bonds of a company have a market price of $415.10, a face value of $1,000 , and a yield to maturity of 9.23 percent. how many years is it until these bonds mature?
Answer #1

Price of bond = Future cash flow / ( 1 + 9.23%)^years

415.10 = 1000/(1+9.23%)^ n

= (1+9.23%)^n = 1000/415.10

= (1.0923)^n = 2.4091

= n = 9.96 years